OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 con.

OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above? (5 points) A) Associative Property of Addition B) Commutative Property of Subtraction C) Addition Property of Equality D) Subtraction Property of Equality

OpenStudy (anonymous):

D) Subtraction Property of Equality

OpenStudy (anonymous):

can you help me with another Josh?

OpenStudy (anonymous):

Step 2; Subtraction Property of Equality. Do you understand why we got that answer?

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

OH! i see it now, thanks soso much!

OpenStudy (anonymous):

cool, no probs liza :)